A ray EF of monochromatic light is incident on the refracting AB of a regular glass prism(refractive index=1.5)at an angle of incidence i=55°.if it emerges through the adjacent face AC..calculate angle of emergence
1
Expert's answer
2020-02-18T09:43:45-0500

As per the question,

Refractive index of the prism "(\mu)=1.5"

Angle of incidence (i)"= 55^\circ"

As per the question, the prism is the regular glass prism, So the prism angle (A)="60^\circ"

Let the angle of deviation of the light "\delta_{min}," and angle of emergence = e

We know that

"\Rightarrow \dfrac{\sin{\dfrac{A+\delta}{2}}}{\sin{\dfrac{A}{2}}}=\mu"

"\Rightarrow \dfrac{\sin{\dfrac{60+\delta}{2}}}{\sin{\dfrac{60}{2}}}=1.5"

"\Rightarrow \dfrac{\sin{\dfrac{A+\delta}{2}}}{\sin{30}}=1.5"


"\Rightarrow 2\sin{\dfrac{A+\delta}{2}}=1.5"

"\Rightarrow \sin{(30+\delta/2)}=\dfrac{3}{4}"

"\Rightarrow 30+\delta/2=48.6^\circ"

"\Rightarrow \delta/2=48.6^\circ-30^\circ=18.6^\circ"

"\Rightarrow \delta=37.2^\circ"

We know that, for the prism,

"i+e=A+\delta"

"\Rightarrow e=A+\delta-i"

"\Rightarrow e=60+37.2-55=42.2^\circ"

Hence, the angle of emergent will be "42.2^\circ"